OpenStudy (anonymous):

Help Partial Fractions \[\int\limits_{}^{}\frac{ dx }{ (x+1)(x+2) }\]

zepdrix (zepdrix):

Any confusion on that setup or multiplication?

OpenStudy (anonymous):

no but before we move on I actually meant to put x-1

OpenStudy (anonymous):

just to clear things up

zepdrix (zepdrix):

\[\Large \color{teal}{\frac{1}{(x-1)(x+2)}\quad=\quad \frac{A}{x-1}+\frac{B}{x+2}}\]Oh ok :) I'm gonna erase the other comment then so we don't get confused.\[\Large \color{teal}{1\quad=\quad A(x+2)+B(x-1)}\]

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

To solve for the unknown constants A and B, we have a couple of options: ~We can multiply out the brackets and equate like terms ( which ends up being really messy in a lot of cases, not this one though ). ~Or we can plug in specific x values that will cancel out one of the unknowns, allowing us to solve for the other.

zepdrix (zepdrix):

For example: if we plug in \(\Large x=1\), what does that give us here?\[\Large \color{teal}{1\quad=\quad A(x+2)+B(x-1)}\]

OpenStudy (anonymous):

a=1/3

zepdrix (zepdrix):

It cancelled out the B, allowing us to solve for A. Cool! So what value x-value would allow us to cancel the A, and solve for B? :X

OpenStudy (anonymous):

-2

zepdrix (zepdrix):

sounds good c:

OpenStudy (anonymous):

b=-1/3

zepdrix (zepdrix):

Ok good, so here is what we've determined.\[\Large \frac{A}{x-1}+\frac{B}{x+2}\quad=\quad \frac{(1/3)}{x-1}+\frac{(-1/3)}{x+2}\]It's always ugly writing fractions on fractions -_- lol

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

\[\large \int\limits\frac{1}{(x-1)(x+2)}dx\quad=\quad \int\limits \frac{A}{x-1}+\frac{B}{x+2}\;dx\] \[\large =\quad \int\limits \frac{(1/3)}{x-1}-\frac{(1/3)}{x+2}\;dx\]

zepdrix (zepdrix):

This will be easier to read if we factor a 1/3 out of each term and pull it outside of the integral.\[\Large =\quad \frac{1}{3}\int\limits \frac{1}{x-1}-\frac{1}{x+2}\;dx\]

zepdrix (zepdrix):

Understand how to solve it from here? :o

OpenStudy (anonymous):

so \[(\frac{ 1 }{ 3 })\ln \frac{ x+2 }{ x-1 }\]

OpenStudy (anonymous):

+C

zepdrix (zepdrix):

Hmm I think your log stuff is upside-down, no?

zepdrix (zepdrix):

\[\Large \log(a)-\log(b)\quad=\quad \log\left(\frac{a}{b}\right)\]

OpenStudy (anonymous):

oh x-2 is negative my fault not x-1

OpenStudy (anonymous):

alright thanks i think I can handle this

zepdrix (zepdrix):

yay team \c:/